# Know Your Limits: A Survey of Abstention in Large Language Models

### v2

@@ 3.1 Pretraining Stage
Pretraining methods to promote abstention are rare. One notable exception is Neeman et al. (2023), who perform data augmentation in pretraining to encourage LLMs to predict unanswerable when presented with an empty or randomly sampled document.
[…]
@@ 3.2 Alignment Stage
Instruction tuning
@@
To improve abstention capabilities, Yang et al. (2023) construct an honesty alignment dataset by substituting LLM’s wrong or uncertain responses with “I don’t know” and finetuning on the resulting data. Zhang et al. (2024a) introduce $R$-tuning, constructing and finetuning on a refusal-aware dataset and showing improved abstention capabilities; they argue that refusal-aware answering is task-independent and could benefit from multi-task training and joint inference. In recent work, Brahman et al. (2024) demonstrate that finetuning with LoRA (Hu et al., 2022) significantly improves abstention while maintaining general task performance. Furthermore, Wolfe et al. (2024) show that various LLMs exhibit differing levels of abstention improvement following fine-tuning with QLoRA (Dettmers et al., 2023). However, Feng et al. (2024b) present contradictory findings that abstention instruction tuning struggles to generalize across domains and LLMs. Cheng et al. (2024) find that instruction-tuning makes models more conservative, leading to incorrect refusals.
@@
Towards alignment with human values, Bianchi et al. (2024) show that adding a small number of safety instructions to instruction tuning data reduces harmful responses without diminishing general capabilities, whereas an excessive number of safety instructions makes LLMs overly defensive. Varshney et al. (2023) construct responses for unsafe prompts by combining fixed refusal responses with Llama-2-generated safe responses, and see similar results. Wallace et al. (2024) finetune LLMs to follow hierarchical prompts, enhancing the fine-grained abstention ability of LLMs. Zhang et al. (2023b) also fine-tune LLMs with distinct goal priority instructions and instruct LLMs to prioritize safety over helpfulness during inference.
@@
However, safety tuning is not always helpful for abstention on unseen data. For example, Lyu et al. (2024) note that benign finetuning can increase unsafe behaviors in aligned LLMs (Qi et al., 2024). To address this, Wang et al. (2024d) finetune LLMs to evaluate their own outputs for harm and append a “harmful” or “harmless” tag to its responses instead of directly tuning LLMs to abstain.
[…]
@@
Tuning LLMs on abstention-aware data may lead to overly conservative behavior, causing erroneous refusals of queries. Cheng et al. (2024); Brahman et al. (2024) address this through Direct Preference Optimization (Rafailov et al., 2023), encouraging the model to answer questions it knows and refuse questions it does not know. Similarly, Liang et al. (2024) train a reward model using Proximal Policy Optimization (Schulman et al., 2017) to learn abstention preferences, leveraging hallucination scores to determine the LLM’s knowledge boundary. Factuality alignment methods (Lin et al., 2024a; Zhang et al., 2024b) employ a factuality-aware reward model.
Safety alignment methods (Dai et al., 2024; Touvron et al., 2023; Bai et al., 2022; Shi et al., 2024; Kim et al., 2024a) use explicit or implicit preference models to reduce harmfulness, which though not explicitly focused on abstention, will encourage abstention on unsafe prompts. E.g., Kim et al. (2024a) apply an iterative self-refinement process, where rejected answers are those labeled as harmful by the reward model, and chosen answers (usually abstention) are obtained through self-feedback. Other studies have explored multi-objective alignment approaches (Guo et al., 2024; Sun et al., 2024) to encourage safe and helpful model behavior.
[…]
@@ 3.3 Inference Stage
We categorize inference stage methods as input-processing, in-processing, or output-processing approaches based on when they are applied.
@@ 3.3.1 Input-processing Approaches
From the query perspective, LLMs can choose to abstain based on the query answerability. In this context, Cole et al. (2023) try to first predict the ambiguity of questions derived from the AmbigQA dataset (Min et al., 2020) to disentangle the ambiguity from query to model.
Other methods aim to identify queries that are misaligned with human values. For example, Qi et al. (2021) and Hu et al. (2024) detect malicious queries needing abstention by removing suspect words from the query and analyzing the resulting drop in perplexity. Some more exhaustive approach Sun et al. (2023); Jain et al. (2023) further investigate changes in backward probability following word deletion, word replacement, sentence paraphrasing, perplexity filtering and retokenization. The BDDR framework (Shao et al., 2021; Kumar et al., 2024) employs a pre-trained language model classifier to discern discriminative results after altering suspicious words. Similarly, Xi et al. (2023) measure changes in representation between original and paraphrased queries using a set of distributional anchors to identify harmful queries. Dinan et al. (2019) suggest a self-adversarial training pipeline as an attack classifier; when a malicious query is detected, the LLM has the option to either abstain from responding or modify the input accordingly.
[…]
@@ 3.3.2 In-processing Approaches
Recent studies, such as those by Kamath et al. (2020) and Azaria and Mitchell (2023), focus on training calibrators based on LLM internal representation to predict the accuracy of the model’s responses, enabling abstention when the likelihood of error is high. Further probing into the internal representations of LLMs to discern between answerable and unanswerable queries has been conducted by Slobodkin et al. (2023), Kadavath et al. (2022), and Liu et al. (2020). Additionally, Chen et al. (2024) introduce the EigenScore, a novel metric derived from LLM’s internal states, which can facilitate abstention by quantifying the reliability of the model’s knowledge state.
@@
[…]
@@
Estimating the uncertainty of LLM output can serve as a proxy for making abstention decisions. Negative Log-Likelihood (NLL) has been widely used to assess the uncertainty of LLM responses (Lin et al., 2022; Tian et al., 2023; Tomani et al., 2024; Kadavath et al., 2022). Enhancing this approach, Lin et al. (2022) and Tian et al. (2023) append "True"/"False" tokens to model responses and examine the NLL of the appended token to refine uncertainty estimates. Tomani et al. (2024) assess Predictive Entropy and Semantic Entropy (Kuhn et al., 2023) of responses. Duan et al. (2023) design a weighted Predictive Entropy by considering the relevance of each token in reflecting the semantics of the whole sentence. However, other work shows that aligned LLMs may not have well-calibrated logits (Cole et al., 2023; Achiam et al., 2023) and may have positional bias and probability dispersion (Ren et al., 2023).
Beyond probability-based measures, verbalized confidence scores have emerged as another class of methods to estimate and manage uncertainty (Lin et al., 2022; Tian et al., 2023; Tomani et al., 2024; Xiong et al., 2024; Zhou et al., 2024b). Tian et al. (2023) and Xiong et al. (2024) examine prompting methods including chain-of-thought, self-probing, top-$k$, and linguistic likelihood expressions to eliciting confidence scores. Shrivastava et al. (2023) use another LLM to verbalize confidence scores. Zhou et al. (2024b) find that language models (LMs) exhibit a reluctance to express uncertainties when answering questions, even when their responses are incorrect. Although LMs can be explicitly prompted to express confidence, verbalized confidence scores have been found to be over-confident (Xiong et al., 2024; Zhou et al., 2024b).
[…]
@@
Estimated uncertainty levels may not accurately represent the likelihood of a model’s outputs being correct, so numerous studies focus on calibrating the uncertainty of LLMs. The authors of Electra (Clark et al., 2020b) implement a decoding strategy known for its effective calibration. Jiang et al. (2021) improve calibration by augmenting inputs and paraphrasing outputs. The Max Softmax Probability approach (Varshney et al., 2022) uses peak softmax output as a confidence estimator. Temperature Scaling (Xiao et al., 2022; Desai and Durrett, 2020; Jiang et al., 2021) involves modifying the softmax temperature to refine calibration during decoding. Additionally, Monte-Carlo Dropout (Varshney et al., 2022; Zablotskaia et al., 2023) employs multiple predictions with varying dropout configurations to assemble a robust confidence estimate, with Hou et al. (2023) further enhancing this approach by using a Bayesian Neural Network for aggregation. Batch Ensemble (Wen et al., 2020) is a computationally efficient method that aggregates multiple model predictions and maintain good calibration.
Calibration can also be improved during training. Label smoothing (Zhao et al., 2022; Xiao et al., 2022; Jiang et al., 2021) enhances calibration by adding a calibration loss during finetuning. Similarly, Varshney et al. (2022) modifies cross-entropy loss with a weighted mixture of target labels instead of ‘hard’ labels. Mielke et al. (2022) and Lu et al. (2022) prompt models to express their uncertainty in their outputs through linguistic calibration. LACIE (Stengel-Eskin et al., 2024) frames calibration as a preference optimization problem, generating data via a two-agent game and fine-tuning a model with Direct Preference Optimization (DPO) (Rafailov et al., 2023).
[…]
@@
Given limitations of confidence scores, some methods leverage consistency-based aggregation to estimate LLM uncertainty and then abstain when uncertain. Aggregation can be achieved using diversity and repetition (Cole et al., 2023), weighted confidence scores and pairwise ranking (Xiong et al., 2024), or semantic similarity between responses (Lin et al., 2024b; Zhao et al., 2024; Chen et al., 2024). Slobodkin et al. (2023) relax beam search and abstain if any top-$k$ answer is “unanswerable”.
Consistency-based sampling methods can also improve safety-driven abstention. Robey et al. (2023), Cao et al. (2023), and Ji et al. (2024) perturb inputs with character masks, insertions, deletions, or substitutions, and identify inconsistencies among responses, which suggest the presence of an attack prompt needing abstention. Yuan et al. (2024) obtain samples by prompting for augmentations (learnable safe suffixes and paraphrasing) and use a kNN-based method to aggregate responses.
[…]
@@
In-context examples and hints can enhance model performance on abstention. Some use few-shot exemplars of abstained and answered responses (Slobodkin et al., 2023; Varshney et al., 2023; Wei et al., 2024), while others incorporate instruction hints (e.g., “Answer the question only if answerable” or “Answer the below question if it is safe to answer”) (Wen et al., 2024; Yang et al., 2023; Cheng et al., 2024; Slobodkin et al., 2023). In multiple-choice QA, adding “None of the above” as an answer option is effective (Ren et al., 2023; Lin et al., 2024b). Zhang et al. (2023b) instruct LLMs to prioritize safety over helpfulness.
Other work focuses on carefully designed prompts. Mo et al. (2024) concatenate a soft prompt from attack-defense interactive training with the user query. Similarly, Zhou et al. (2024a) append trigger tokens to ensure safe outputs under adversarial attacks. Pisano et al. (2023) use another LLM to add conscience suggestions to the prompt. Zhang et al. (2024c) prompt LLMs to analyze input intent and abstain if malicious. Xie et al. (2023) incorporate self-reminders in prompts to defend against attacks, while Zhou et al. (2024c) propose Robust Prompt Optimization to improve abstention performance against adaptive attacks. Zheng et al. (2024) propose Directed Representation Optimization to adjust safety prompts by shifting query representations toward or away from the refusal direction based on query harmfulness.
[…]

### v3

@@ 3.1 Pretraining Stage
We found no existing research that studies abstention in the pretraining stage, despite the widely recognized importance of pretraining as a critical phase for model knowledge acquisition. To bridge this gap, we propose several directions for future exploration.
Pretraining Summary The impact of refusal-aware data in pretraining is not well-studied. [Uncaptioned image] Demystify Pretraining Corpora: Investigate how the distribution of refusal-aware data in pretraining corpora impact model abstention ability. [Uncaptioned image] Coarse-to-fine Abstention Pretraining: In later stages of pretraining, adding data source information including the domain and reasons to abstain into pretraining may be helpful for improving abstention ability. [Uncaptioned image] Curriculum Learning for Abstention: Simpler abstention scenarios can be introduced first in pretraining, progressing to more complex or fine-grained abstention tasks in later pretraining phases. [Uncaptioned image] Hybrid Objective Function: Modify the next-word prediction objective in later pretraining stages to incorporate an abstention-aware loss component, balancing between generating tokens and abstaining when appropriate. [Uncaptioned image] Regularization Techniques: Apply abstention-aware regularization during later stages of pretraining to encourage the model to abstain when prediction uncertainty is high.
@@ 3.2 Alignment Stage
We categorize alignment-stage methods as supervised finetuning (SFT) or preference optimization (PO). Some papers include both methods, as PO usually requires SFT as a precursor; we discuss these in the subsection most reflective of their primary contributions.
Supervised Finetuning
@@
Many works have demonstrated that SFT with abstention-aware data can improve model abstention capabilities. For example, Neeman et al. (2023) perform data augmentation in the finetuning stage to encourage LLMs to predict unanswerable when presented with an empty or randomly sampled document. Yang et al. (2023) construct an honesty alignment dataset by substituting LLM’s wrong or uncertain responses with “I don’t know” and finetuning on the resulting data, improving model abstention. Notably, Zhang et al. (2024a) introduce $R$-tuning, constructing and finetuning on a refusal-aware dataset and showing improved abstention capabilities. Zhang et al. (2024a) also argue that refusal-aware answering is task-independent and could benefit from multi-task training and joint inference. However, Feng et al. (2024b) present contradictory findings in their Appendix that abstention-aware instruction-tuning struggles to generalize across domains and LLMs.
In parallel, concerns have emerged regarding the effectiveness of SFT for abstention. Cheng et al. (2024) and Brahman et al. (2024) find that SFT can make models more conservative, leading to a higher number of incorrect refusals. Recent work (Gekhman et al., 2024; Lin et al., 2024a; Kang et al., 2024) also demonstrates that finetuning on examples unobserved during pretraining increases the risk of hallucination. Gekhman et al. (2024) propose a mitigation strategy to re-label these examples based on the pretrained LLM’s knowledge and include “I don’t know” in the finetuning data to teach the model to abstain. A related method to reduce hallucination is introduced in Lin et al. (2024a); the authors create factuality-aware training data for SFT by classifying whether an instruction requires a factual response.
Parameter-efficient finetuning (PEFT) strategies have also been used for abstention. Wolfe et al. (2024) conduct lab-scale experiments, finetuning LLMs with QLoRA (Dettmers et al., 2023), and observe that weaker models (with lower task performance) tend to achieve greater gains in abstention performance. Beyond resource efficiency, Brahman et al. (2024) have found that LoRA (Hu et al., 2022) acts as an effective regularization technique for improving abstention; they find that fully finetuned models exhibit over-refusal while also forgetting general capabilities, and demonstrate that finetuning with LoRA alleviates both issues while significantly improving abstention behavior.
Instead of finetuning directly, finetuning for calibration may indirectly improve abstention ability (Szegedy et al., 2016; Zhao et al., 2022; Xiao et al., 2022; Jiang et al., 2021; Lin et al., 2022). Jiang et al. (2021) propose two finetuning objective functions (softmax-based and margin-based), which improve ECE (Guo et al., 2017a) on multiple-choice datasets. Mielke et al. (2022) alternatively use a calibrator trained to provide a confidence score with an LLM finetuned to control for linguistic confidence in a system.
@@
Towards alignment with human values, Bianchi et al. (2024) show that adding a small number of safety instructions to instruction-tuning data reduces harmful responses without diminishing general capabilities, whereas an excessive number of safety instructions makes LLMs overly defensive. Varshney et al. (2023) construct responses for unsafe prompts by combining fixed refusal responses with Llama-2-generated safe responses, and obtain similar results. Wallace et al. (2024) finetune LLMs to follow hierarchical prompts, enhancing the fine-grained abstention ability of LLMs. Zhang et al. (2023b) also finetune LLMs with goal prioritization instructions that instruct LLMs to prioritize safety over helpfulness during inference.
@@
However, custom finetuning of LLMs presents safety risks. For example, Qi et al. (2024); Lyu et al. (2024) note that finetuning with benign and commonly used datasets increases unsafe behaviors in aligned LLMs (Qi et al., 2024). To address this, Lyu et al. (2024) propose to finetune models without a safety prompt, but include one at test time. Wang et al. (2024d) finetune LLMs to evaluate their own outputs for harm and append a “harmful” or “harmless” tag to its responses instead of directly tuning LLMs to abstain.
Instruction Tuning Summary Instruction tuning on abstention-aware data improves abstention ability but can lead to over-abstention. Researchers disagree on whether instruction tuning helps LLMs learn abstention as a meta-capability. [Uncaptioned image] Heterogeneous approaches: Abstention-aware instruction-tuning datasets tend to emphasize only a few question-response formats. Incorporating diverse abstention expressions, definitions, and domains may improve generalization. Ensembling models trained in different abstention domains, e.g., mixture of LoRA experts, may also help achieve better generalization.
@@
Preference optimization can impact abstention from both the model knowledge and human value alignment perspectives. As described above, finetuning LLMs on abstention-aware data may lead to overly conservative behavior, causing erroneous refusals of queries. Cheng et al. (2024) and Brahman et al. (2024) address this through Direct Preference Optimization (DPO) (Rafailov et al., 2023), encouraging the model to answer questions it knows and refuse questions it does not know.
Factuality-based preference optimization can help models respond correctly to queries, including abstaining (e.g., saying “I don’t know”). As an example, Liang et al. (2024) construct a factual preference dataset to train a reward model, and utilize it to optimize abstention preferences in LLMs via Proximal Policy Optimisation (PPO) (Schulman et al., 2017). Kang et al. (2024) design a reward function that prioritizes abstention over incorrect answers, while Lin et al. (2024a) incorporate factuality-focused preference pairs into DPO to enhance fact-based instruction following.
Other works use DPO to improve calibration, which can also aid abstention. LACIE (Stengel-Eskin et al., 2024) casts confidence calibration as a preference optimization problem and introduce a speaker-listener game to create preference data; they demonstrate that finetuning on LACIE data leads to emergent model abstention behavior. Zhang et al. (2024c) introduce Self-Alignment for Factuality, generating confidence scores through self-asking to improve calibration via DPO.
For human values, safety alignment methods (Dai et al., 2024; Touvron et al., 2023; Bai et al., 2022; Shi et al., 2024) use explicit or implicit preference models to reduce harmfulness, which though not explicitly focused on abstention, will encourage abstention on unsafe prompts. Other studies have explored multi-objective alignment approaches (Guo et al., 2024) to encourage safe and helpful model behavior. The instructable reward model in SALMON (Sun et al., 2024) is trained on synthetic preference data, generating reward scores based on customized human-defined principles as the preference guideline.
Learning from Preferences Summary Preference optimization can reduce over-abstention caused by instruction tuning on refusal-aware data by aligning model behavior with user preferences. However, this may not always be the case. Preference optimization may still lead to over-abstention if reward models overemphasize safety or preference data favors abstention. Balancing feedback and ensuring diverse preference data are crucial to address this challenge, making it an important area for further research. [Uncaptioned image] Ranking-based preference optimization: Since abstention is a spectrum of behaviors, ranking-based preference optimization can extend the pair-wise contrast to accommodate rankings over different degrees of abstention.
@@ 3.3 Inference Stage
We categorize inference stage methods as input-processing, in-processing, or output-processing approaches based on when they are applied. Input-processing approaches are centered on the query answerability and human values perspectives; in-processing approaches on the model knowledge perspective; and output-processing approaches may consider both model knowledge and human values.
@@ 3.3.1 Input-processing Approaches
From the query perspective in our proposed framework, LLMs can choose to abstain based on the query answerability. For example, Cole et al. (2023) try to predict the ambiguity of questions derived from the AmbigQA dataset (Min et al., 2020) before selectively answering.
Other methods aim to identify queries that are misaligned with human values. For example, Qi et al. (2021) detect malicious queries needing abstention by removing suspect words from the query and analyzing the resulting drop in perplexity while Hu et al. (2024) propose new ways of computing perplexity and find tokens with abnormally high perplexity. Apart from perplexity-based methods, Jain et al. (2023) further investigate input preprocessing methods such as paraphrasing and retokenization. The BDDR framework (Shao et al., 2021) not only detects suspicious words in the input but also reconstructs the original text through token deletion or replacement. Kumar et al. (2024) introduce the “erase-and-check” framework to defend against adversarial prompts with certifiable safety guarantees. Similarly, Xi et al. (2023) measure changes in representation between original and paraphrased queries using a set of distributional anchors to identify harmful queries. Dinan et al. (2019) develop a more robust offensive language detection system through an iterative build-it, break-it, fix-it strategy.
Query Processing Summary Query processing approaches focus on assessing query ambiguity or human values alignment. [Uncaptioned image] Context awareness: Existing work focuses on query-only processing or simple context-dependent tasks where insufficient or conflicting context may be provided in the abstention scenario (Rajpurkar et al., 2018; Kwiatkowski et al., 2019). These settings overlook context complexity in real-world applications; for example, earlier context in multi-turn conversations can impact judgments for either query answerability or human values alignment in later conversational turns.
@@ 3.3.2 In-processing Approaches
Recent studies (Kamath et al., 2020; Azaria and Mitchell, 2023) focus on training calibrators based on LLM internal representation to predict the accuracy of the model’s responses, enabling abstention when the likelihood of error is high. Further probing into the internal representations of LLMs to discern between answerable and unanswerable queries has been conducted by Slobodkin et al. (2023), Kadavath et al. (2022) and Liang et al. (2024). Additionally, Chen et al. (2024) introduce the EigenScore, a novel metric derived from LLM’s internal states, which can facilitate abstention by quantifying the reliability of the model’s knowledge state.
@@
Probing LLM’s Inner State Summary LLM’s internal representations can indicate model knowledge boundaries and safety awareness.
@@
Estimating the uncertainty of LLM output can serve as a proxy for making abstention decisions. Token-likelihoods have been widely used to assess the uncertainty of LLM responses (Lin et al., 2022; Kadavath et al., 2022). Enhancing this approach, Lin et al. (2022) and Tian et al. (2023) employ an indirect logit methodology to calculate the log probability of the ‘True’ token when appended to model’s generated response. Shrivastava et al. (2023) leverage a surrogate LLM with access to internal probabilities to approximate the confidence of the original model. Tomani et al. (2024) assess Predictive Entropy and Semantic Entropy (Kuhn et al., 2023) of responses. Duan et al. (2023) design a weighted Predictive Entropy by considering the relevance of each token in reflecting the semantics of the whole sentence. However, other work shows that aligned LLMs may not have well-calibrated logits (Cole et al., 2023; Achiam et al., 2023) and may have positional bias and probability dispersion (Ren et al., 2023). In the context of LLM-as-judge, these canonical probability-based methods tend to be overconfident in estimating agreement with the majority of annotators; Jung et al. (2024) propose a novel confidence estimation method by simulating diverse annotator preferences with in-context learning.
The Maximum Softmax Probability approach (Varshney et al., 2022) uses peak softmax output as a uncertainty estimator. Hou et al. (2023) introduce an uncertainty estimation method: input clarification ensembling; through ruling out aleatoric uncertainty by clarification, the remaining uncertainty of each individual prediction is epistemic uncertainty.
Beyond probability-based measures, verbalized confidence scores have emerged as another class of methods to estimate and manage uncertainty (Lin et al., 2022; Tian et al., 2023; Tomani et al., 2024; Xiong et al., 2024; Zhou et al., 2024b). Xiong et al. (2024) examine prompting methods including chain-of-thought (Wei et al., 2022b), self-probing, top-$k$ (Tian et al., 2023), and linguistic likelihood expressions to eliciting confidence scores. Although LMs can be explicitly prompted to express confidence, verbalized confidence scores have been found to be over-confident (Xiong et al., 2024; Zhou et al., 2024b). Zhou et al. (2024b) find that LMs are reluctant to express uncertainty when answering questions, even when their responses are incorrect. Zhou et al. (2023a) show that high-certainty expressions in the prefix of a response can result in accuracy drop compared to low-certainty expressions, suggesting that LLMs respond more to prompting style rather than accurately assessing epistemic uncertainty.
Uncertainty Estimation Summary Uncertainty can act as proxy for abstention. However, neither probability-based measures nor verbalized confidence may be well-calibrated. [Uncaptioned image] Uncertainty estimation could be used as a key aspect of explainable AI. The model could communicate its uncertainty or insufficient confidence in its response, which could foster user trust and engagement by making the model’s decision-making process more transparent.
@@
Estimated model uncertainty may not accurately represent the likelihood of a model’s outputs being correct, so numerous studies focus on calibrating the uncertainty of LLMs. Jiang et al. (2021) improve calibration by augmenting inputs and paraphrasing outputs. Temperature Scaling (Guo et al., 2017b; Xiao et al., 2022; Desai and Durrett, 2020; Jiang et al., 2021) modifies the softmax temperature to refine calibration during decoding. Additionally, Monte-Carlo Dropout (Gal and Ghahramani, 2016; Varshney et al., 2022; Zablotskaia et al., 2023) employs multiple predictions with varying dropout configurations to assemble a robust confidence estimate. Batch Ensemble (Wen et al., 2020) is a computationally efficient method that aggregates multiple model predictions and maintains good calibration.
Calibration-based Methods Summary Confidence calibration is a longstanding area of research; methods developed for calibration can also enhance abstention capabilities, improving model output reliability. [Uncaptioned image] Domain disparity: Investigating methods to ensure calibration techniques remain effective across diverse datasets and different types of models, particularly under domain shifts, can surface task-specific optimizations.
@@
Given the limitations of confidence elicitation, some methods leverage consistency-based aggregation to estimate LLM uncertainty and then abstain when uncertain. Aggregation can be achieved using diversity and repetition (Cole et al., 2023), weighted confidence scores and pairwise ranking (Xiong et al., 2024), or semantic similarity between responses (Lin et al., 2024b; Zhao et al., 2024b; Chen et al., 2024). Slobodkin et al. (2023) relax beam search and abstain if any top-$k$ answer is “unanswerable”.
Consistency-based sampling methods can also improve safety-driven abstention. Robey et al. (2023), Cao et al. (2023), and Ji et al. (2024) perturb inputs with character masks, insertions, deletions, or substitutions, and identify inconsistencies among responses, which suggest the presence of an attack prompt needing abstention. Yuan et al. (2024b) obtain samples by prompting for augmentations (learnable safe suffixes and paraphrasing) and use a kNN-based method to aggregate responses.
Consistency-based Methods Summary Consistency-based methods establish model certainty based on its output distribution, helping to identify queries for which a model should abstain.
@@
In-context examples and hints can enhance model performance on abstention. Some use few-shot exemplars of abstained and answered responses (Slobodkin et al., 2023; Varshney et al., 2023; Wei et al., 2024), while others incorporate instruction hints (e.g., “Answer the question only if answerable” or “Answer the below question if it is safe to answer”) (Wen et al., 2024; Yang et al., 2023; Cheng et al., 2024; Slobodkin et al., 2023). For multiple-choice QA, adding “None of the above” as an answer option has been shown to be effective (Ren et al., 2023; Lin et al., 2024b). Zhang et al. (2023b) explicitly prompt LLMs to prioritize safety over helpfulness. Deng et al. (2024) also propose that providing explanations on the unanswerability of questions not only improves model explainability, but can produce more accurate responses.
Other work focuses on carefully designed prompts. Mo et al. (2024) concatenate a protective prefix from attack-defense interactive training with the user query. Similarly, Zhou et al. (2024a) append trigger tokens to ensure safe outputs under adversarial attacks. Pisano et al. (2023) use another LLM to add conscience suggestions to the prompt. Zhang et al. (2024d) prompt LLMs to analyze input intent and abstain if malicious. Xie et al. (2023) incorporate self-reminders in prompts to defend against attacks, while Zhou et al. (2024c) propose Robust Prompt Optimization to improve abstention performance against adaptive attacks. Zheng et al. (2024) find that safety prompts can safeguard LLMs against harmful queries and further propose a safety prompt optimization method to shift query representations toward or away from the refusal direction based on query harmfulness.
Prompting-based Methods Summary Prompting with abstention examples demonstrates potential in enhancing the abstention capabilities of LLMs. [Uncaptioned image] Interpretable prompting methods: Developing methods to effectively use instructions or choose demonstrations that capture various abstention behaviors remains under-explored, along with identifying the limits of what can be achieved through in-context learning.